To save money for a new house, you want to begin contributing money to a savings account. Your plan



is to make equal contributions to the account for eight years. Each contribution will be R18 490,00. The



contributions will come at the end of every three month period, starting three months from now. The savings



account pays 9% interest per annum, compounded quarterly. How much money do you expect to have in



the account at the end of the eight years

1
Expert's answer
2022-04-01T13:47:54-0400

"FV=pmt[\frac{(1+\frac{r}{m})^{nm}-1}{(\frac{r}{m})}]"

Pmt =R18490

R=9%

N=8 years

M=4


"FV_8=R18490[\frac{(1+\frac{9\%}{4})^{8×4}-1}{(\frac{9\%}{4})}]"

"FV_8=R853090"

The amount in the account tat the end of eight (8) years is R853090